OpenStudy (anonymous):

If an automobile air bag has a volume of 11.4L , what mass of NaN3 (in g) is required to fully inflate the air bag upon impact? Assume STP conditions.

OpenStudy (aaronq):

PV=nRT sub in, n=m/M, solve for m

OpenStudy (theeric):

I agree with aaronq. Standard temperature and pressure (STP) is a temperature value and pressure value commonly used in experiments. I get my values for them from http://en.wikipedia.org/wiki/Standard_conditions_for_temperature_and_pressure . \(P~V=n~R~T\) is the equation that describes the relationships between the following properties when the gas behaves nearly ideally. Gasses will do that at STP, a.k.a. standard temperature and pressure. \(P\) is pressure - here, \(100\ [kPa]\) \(V\) is volume - here, \(11.4\ [L]\) \(n\) is number of molecules expressed in moles, which is the total mass (combined mass of all molecules), \(m\), divided by the mass of just one molecule (the sum of the atomic masses), \(M\). \(R\) is the ideal gas constant, \(8.314\ \left[\dfrac{L\ kPa}{\ ^\circ K\ mol}\right] \). \(T\) is the temperature - here, \(273.15\ [\ ^\circ K]\) To find \(M\), you'll need to add the atomic mass of NA with the atomic mass of three N 's. So, use those values unless your class gave you others. Do as aaronq suggested, and you should come to the correct answer! Feel free to ask more questions on this.